# Approximate Degrees of Multisymmetric Properties with Application to Quantum Claw Detection

### v1

@@ Abstract
The claw problem is central in the fields of theoretical computer science as well as cryptography. The optimal quantum query complexity of the problem is known to be $\Omega\left(\sqrt{G}+(FG)^{1/3}\right)$ for input functions $f\colon[F]\to Z$ and $g\colon[G]\to Z$. However, the lower bound was proved when the range $Z$ is sufficiently large (i.e., $\lvert Z\rvert=\Omega(FG)$). The current paper proves the lower bound holds even for every smaller range $Z$ with $\lvert Z\rvert\geq F+G$. This implies that $\Omega\left(\sqrt{G}+(FG)^{1/3}\right)$ is tight for every such range. In addition, the lower bound $\Omega\left(\sqrt{G}+F^{1/3}G^{1/6}M^{1/6}\right)$ is provided for even smaller range $Z=[M]$ with every $M\in[2,F+G]$ by reducing the claw problem for $\lvert Z\rvert=F+G$. The proof technique is general enough to apply to any $k$-symmetric property (e.g., the $k$-claw problem), i.e., the Boolean function $\Phi$ on the set of $k$ functions with different-size domains and a common range such that $\Phi$ is invariant under the permutations over each domain and the permutations over the range. More concretely, it generalizes Ambainis’s argument [Theory of Computing, 1(1):37-46] to the multiple-function case by using the notion of multisymmetric polynomials.
@@ 1.1 Background
Due to their simplicity and broad applicability, the collision detection/finding problem and its variant, the claw detection/finding problem, are central problems in the fields of theoretical computer science (e.g.,[14, 20, 2, 6, 10, 31]) as well as cryptography (e.g.,[13, 23, 27, 25, 17]). For given function $f$, the goal of the collision detection is to detect the existence of a pair of elements $(x,y)$ such that $f(x)=f(y)$ (since the collision detection without any promise on the input $f$ is often called the element distinctness, we may use the latter term when it is appropriate). As a natural variant of the collision detection, we can define the claw detection for two functions $f,g$ as the problem of detecting the existence of a pair $(x,y)$ such that $f(x)=g(y)$. When we are required to find a collision (a claw) if it exists, we refer to the corresponding “search” problem as the collision finding (the claw finding, respectively) in this paper.
@@
For large domain(s), these problems are generally hard (i.e., time-consuming) to solve. This hardness forms the basis of the security of various cryptographic applications. This fact in turn makes these problems appealing as targets with which to investigate quantum computational power: If one finds quantum algorithms that solve the problems very efficiently, this exhibits solid quantum advantages concerning practically essential problems; if one proves the hardness of the problems even against quantum computation, this gives theoretical evidence that the cryptographic applications are secure against quantum computers.
@@
Interestingly, the actual situation is somewhere in between. For these problems, quantum speed-ups are possible but at most polynomials. This has motivated the invention of various important techniques for deriving quantum upper bounds (i.e., quantum algorithms) and lower bounds.
@@
For the element distinctness, Buhrman et al. [14] proved the upper bound $O(n^{3/4})$ and the lower bound $\Omega(\sqrt{n})$ on the quantum query complexity for $f\colon[n]\to Z$ for a finite set $Z$. Aaronson and Shi [2] improved the lower bound to $\Omega(n^{2/3})$ by introducing a novel technique to the polynomial method to bound the degrees of multi-variate polynomials that approximate symmetric Boolean functions, where $[n]$ denotes the set $\{1,\dots,n\}$. Ambainis [6] then provided the breakthrough upper bound $O(n^{2/3})$ by developing a general framework of quantum search based on quantum walk. His work was followed by many generalizations and variants that form one of the mainstreams of quantum algorithms (e.g.,[29, 26, 8, 22]).
@@
As for the claw detection for $f\colon[F]\to Z$ and $g\colon[G]\to Z$, where $F$ and $G$ are positive integers such that $F\leq G$ and $Z$ is a finite set, Buhrman et al. [14] proved that the quantum query complexity is $\Omega(\sqrt{G})$ and $O(\sqrt{G}+F^{1/2}G^{1/4})$. Zhang [32] then proved the lower bound $\Omega(\sqrt{G}+(FG)^{1/3})$ on the quantum query complexity by providing the reduction from the collision detection with promised that the input function with domain $[FG]$ is either one-to-one or two-to-one and using the lower bound $\Omega((FG)^{1/3})$ [2] for that promised case. Tani [30] then proved the matching upper bound $O(\sqrt{G}+(FG)^{1/3})$ by running a quantum-walk-based search on a particular graph product of two Johnson graphs corresponding to the two input functions.
@@
These tight bounds depend only on the domain size(s) of the input function(s). This means that these bounds are tight if the function range is sufficiently large. Indeed, the lower bound $\Omega(n^{2/3})$ of the element distinctness was originally proved when the range size is at least $1.5n$ [2]. In some cases of smaller range size, however, it is still non-trivial to detect a collision (e.g., a collision for a given function $f\colon[n]\to[n]$ since $f$ can be one-to-one). Moreover, in some applications, such as cryptographic ones, the range size is often known and can be a relatively small value. Ambainis [4] and Kutin [24] independently proved the same lower bound $\Omega(n^{2/3})$ holds even for the range of size $n$. In particular, Ambainis’s argument is so general that it can be applied to any function $\Phi$ from a set of functions $f$ to $\{0,1\}$ if $\Phi$ is invariant under the permutations $\pi$ and $\sigma$ over the domain and range, respectively, of $f$, i.e., $\Phi(f)=\Phi(\sigma f\pi)$ (e.g., if $\Phi$ is the function that tells the existence of a collision for $f$, then it is easy to see that $\Phi(f)=\Phi(\sigma f\pi)$).
@@
As for the claw detection, Zhang’s lower bound holds if the range size of input functions is $\Omega(FG)$. This condition seems very strong, considering that it is still non-trivial to detect a claw for a given pair of functions $f\colon[F]\to[F+G]$ and $g\colon[G]\to[F+G]$ with $F\leq G$, since the images of $f$ and $g$ can be disjoint. Therefore, it is natural to ask whether the lower bound $\Omega(\sqrt{G}+(FG)^{1/3})$ still holds for a smaller range, such as $[F+G]$, and, more ambitiously, whether a multi-function version of Ambainis’s argument exists. From a practical point of view, relaxing the condition on the function range for tight bounds could contribute to more accurate security analysis of post-quantum ciphers, such as those based on supersingular elliptic curve isogenies [21], one form (SIKE) of which had been considered until the final round of the post-quantum standardization at NIST.
@@ 1.2 Polynomial Method
There are two primary lower-bound techniques for quantum query complexity: the polynomial method [9] and the adversary method [3, 18]. Although the polynomial method does not necessarily give optimal lower bounds [5], it often gives non-trivial lower bounds that are not known how to obtain using other lower-bound techniques. The highlight of the method is the claim that the minimum degree of a polynomial that approximates a Boolean function $\Phi$ is at most twice the number of quantum queries required to compute $\Phi$. This allows us to use the results of polynomial approximation theory to obtain lower-bounds on the quantum query complexity of $\Phi$. More concretely, let $f$ be any function from $[n]$ to $[m]$. Then, $f$ can be identified with $nm$ Boolean variables $\{x_{ij}:i\in\left[n\right],j\in\left[m\right]\}$ such that $x_{ij}=1$ if and only if $f\left(i\right)=j$. Let $\Phi$ be a Boolean function that tells whether input $f$ has a certain property (e.g., the property that $f$ has at least one collision). Then, $\Phi$ is a function in the variables $\{x_{ij}\}$. The polynomial method claims that the minimum degree of a polynomial that approximates $\Phi$ within error $\epsilon$ is at most twice the number of quantum queries required to compute $\Phi$ with error probability at most $\epsilon$. For $m\geq n$, suppose that $\Phi^{\prime}$ is the restriction of $\Phi$ to the case where the range of input function is $\left[n\right]$. Then, it is easy to see that if $d$ is the minimum degree of a polynomial that approximates $\Phi$, then the minimum degree of a polynomial that approximates $\Phi^{\prime}$ is at most $d$ (since $\Phi^{\prime}$ is obtained by setting $x_{ij}=0$ for all $i\in[n]$ and $j\in\left[m\right]\setminus\left[n\right]$ in $\Phi$). Ambainis [4] proved that the converse is also true if $\Phi$ is a symmetric property, i.e., a Boolean function invariant under the permutations over domain $[n]$ and range $\left[m\right]$, respectively, of $f$. For the property $\Phi$ that input $f\colon[n]\to[3n/2]$ has at least one collision, the minimum approximate degree of $\Phi$ is known to be $\Omega\left(n^{2/3}\right)$. Thus, the polynomial method together with Ambainis’s argument implies that the quantum query complexity of the element distinctness problem is $\Omega\left(n^{2/3}\right)$ even if the function range is $\left[n\right]$. This relaxes the condition (i.e., $m\geq 1.5n$ ) on the range size of input functions required in [2] to obtain the optimal lower bound.
@@ 1.3 Our Contribution
Our main contribution is to prove a multi-function version of Ambainis’s argument. As an application of our argument to two functions $f\colon[F]\to[M]$ and $g\colon[G]\to[M]$, we show that the optimal quantum lower bound for claw detection proved for range size $FG$ still holds for smaller range size $F+G$. Moreover, by using this new lower bound, we obtain a lower bound for an even smaller range $[M]$ with $2\leq\left|M\right|<F+G$. In the following, we summarize our results more concretely. To simplify the notation, we represent by $\mathcal{R}^{\mathcal{D}}$ the set of functions from $\mathcal{D}$ to $\mathcal{R}$ for any ordered sets $\mathcal{D}$ and $\mathcal{R}$. For instance, if $f$ is a function from $[F]$ to $[M]$, then $f\in[M]^{[F]}$.
@@ 1.3.1 Approximate Degree of Multisymmetric Properties
For $F,G,M\in\mathbb{N}$ such that $M\geq F+G$, let $\Phi\colon[M]^{[F]}\times[M]^{[G]}\to\{0,1\}$ be a bisymmetric Boolean function of a pair of functions $f\colon[F]\to[M]$ and $g\colon[G]\to[M]$. In addition, for every $M^{\prime}$ with $F+G\leq M^{\prime}<M$, $\Phi^{\prime}\colon[M^{\prime}]^{[F]}\times[M^{\prime}]^{[G]}\to\{0,1\}$ be the restriction of $\Phi$ to the case where the input functions are $f\colon[F]\to[M^{\prime}]$ and $g\colon[G]\to[M^{\prime}]$. Then, for $20<\epsilon<1/$ 2, the minimum degree of a polynomial that $\epsilon$-approximates $\Phi$ is equal to the minimum degree of a polynomial that $\epsilon$-approximates $\Phi^{\prime}$.
@@
The definition of bisymmetric properties can be generalized in a straightforward way: Let $\Phi\colon\mathcal{F}_{1}\times\cdots\times\mathcal{F}_{k}\to\{0,1\}$ be a Boolean function over the set of $k$-tuples of functions in $\mathcal{F}_{1}\times\cdots\times\mathcal{F}_{k}$, where $\mathcal{F}_{\ell}\colon[F_{\ell}]\to[M]$ for $F_{\ell},M\in\mathbb{Z}_{+}$. As in the case of two functions, we can see that $\Phi$ depends on Boolean variables $x^{(\ell)}_{ij}\in\{0,1\}$ for $(\ell,i,j)\in[k]\times[F_{\ell}]\times[M]$ such that $x^{(\ell)}_{ij}=1$ if and only if $f_{\ell}\left(i\right)=j$. We say that $\Phi$ is $k$-symmetric, if $\Phi(f_{1},\dots,f_{k})$ is invariant under the permutations over $[F_{\ell}]$ for each $\ell\in[k]$, respectively, and the common range $[M]$ of $f_{\ell}$’s, i.e.,
@@
for every permutation $\pi_{F_{i}}$ over $[F_{i}]$ for $i\in[k]$ and every permutation $\sigma$ over $[M]$. Then, we have a generalization of Theorem 1.
@@
Let $k\geq 2$. For $F_{\ell}\ (\ell\in[k])$ and $M\in\mathbb{N}$ such that $M\geq F_{1}+\dots+F_{k}$, $\Phi:\mathcal{F}_{1}\times\dots\times\mathcal{F}_{k}\to\{0,1\}$ be a $k$-symmetric Boolean function over the set of $k$-tuples of functions $(f_{1},\dots,f_{k})\in\mathcal{F}_{1}\times\dots\times\mathcal{F}_{k}$, where $\mathcal{F}_{\ell}$ denotes $[M]^{[F_{\ell}]}$ for each $\ell$. In addition, for every ${M^{\prime}}$ such that $F_{1}+\dots+F_{k}\leq{M^{\prime}}<M$, let $\Phi^{\prime}:\mathcal{F}^{\prime}_{1}\times\dots\times\mathcal{F}^{\prime}_{k}\to\{0,1\}$ be the restriction of $\Phi$ to the domain $[{M^{\prime}}]^{[F_{1}]}\times\dots\times[{M^{\prime}}]^{[F_{k}]}$, i.e., the $k$-symmetric Boolean function over the set of $k$-tuples of functions $(f^{\prime}_{1},\dots,f^{\prime}_{k})\in\mathcal{F}^{\prime}_{1}\times\dots\times\mathcal{F}^{\prime}_{k}$, where $\mathcal{F}^{\prime}_{\ell}$ denotes $[{M^{\prime}}]^{[F_{\ell}]}$ for each $\ell\in[k]$, such that $\Phi^{\prime}(f^{\prime}_{1},\dots,f^{\prime}_{k})=\Phi(f_{1},\dots,f_{k})$ for all functions $f^{\prime}_{1},\dots,f^{\prime}_{k},f_{1},\dots,f_{k}$ satisfying $f^{\prime}_{\ell}(x)=f_{\ell}(x)$ for all $x\in[F_{\ell}]$ and all $\ell\in[k]$.
@@
With the promise that the union of the images of input functions, i.e., $\bigcup_{\ell\in[k]}{f_{\ell}([F_{\ell}])}$, is upper-bounded in size by $\mathfrak{U}$, Theorem 2 also holds for every $M^{\prime}$ with $\mathfrak{U}\leq M^{\prime}<M$ with a slight modification to the proof. Slightly general versions of the above theorems are Theorems 12 and 15 proved in Sec. missing  3.

### v2

@@ Abstract
The claw problem is central in the fields of theoretical computer science as well as cryptography. The goal of the problem is to detect the existence of a pair of inputs $x$ and $y$ (or find such a pair if they exist) such that $f(x)=g(y)$ for a given pair of funcitons $f$ and $g$ as oracles (i.e., black boxes). The optimal quantum query complexity of the problem is known to be $\Omega\left(\sqrt{G}+(FG)^{1/3}\right)$ for input functions $f\colon[F]\to Z$ and $g\colon[G]\to Z$. However, the lower bound was proved when the range $Z$ is sufficiently large (i.e., $\lvert Z\rvert=\Omega(FG)$). The current paper proves the lower bound holds even for every smaller range $Z$ with $\lvert Z\rvert\geq F+G$. This implies that $\Omega\left(\sqrt{G}+(FG)^{1/3}\right)$ is tight for every such range. In addition, the lower bound $\Omega\left(\sqrt{G}+F^{1/3}G^{1/6}M^{1/6}\right)$ is provided for even smaller range $Z=[M]$ with every $M\in[2,F+G]$ by reducing the claw problem for $\lvert Z\rvert=F+G$. The proof technique is general enough to apply to any $k$-symmetric property (e.g., the $k$-claw problem), i.e., the Boolean function $\Phi$ on the set of $k$ functions with different-size domains and a common range such that $\Phi$ is invariant under the permutations over each domain and the permutations over the range. More concretely, it generalizes Ambainis’s argument [Theory of Computing, 1(1):37-46] to the multiple-function case by using the notion of multisymmetric polynomials.
@@ 1.1 Background
Due to their simplicity and broad applicability, the collision detection/finding problem and its variant, the claw detection/finding problem, are central problems in the fields of theoretical computer science (e.g.,[14, 21, 2, 6, 10, 32]) as well as cryptography (e.g.,[13, 24, 28, 26, 18]). For a given function $f$, the goal of the collision detection is to detect the existence of a pair of elements $(x,y)$ such that $f(x)=f(y)$ (Since the collision detection without any promise on the input $f$ is often called element distinctness, we may use the latter term when appropriate.) As a natural variant of the collision detection, we can define the claw detection for two functions $f,g$ as the problem of detecting the existence of a pair $(x,y)$ such that $f(x)=g(y)$. When we are required to find a collision (a claw) if it exists, we refer to the corresponding “search” problem as the collision finding (the claw finding, respectively) in this paper.
@@
These problems are classically hard (i.e., time-consuming) to solve for large domains. This hardness forms the basis of the security of various cryptographic applications. This fact in turn makes these problems appealing as targets with which to investigate quantum computational power: If one finds quantum algorithms that solve the problems very efficiently, this exhibits solid quantum advantages concerning practically essential problems; if one proves that the problem is very hard even against quantum computation, this gives theoretical evidence that the cryptographic applications are secure against quantum computers.
@@
Interestingly, the actual situation is somewhere in between. For these problems, quantum speed-ups are possible but at most polynomial. This has motivated the invention of various essential techniques for improving quantum upper bounds (i.e., quantum algorithms) and lower bounds.
@@
For element distinctness, Buhrman et al. [14] proved the upper bound $O(n^{3/4})$ and the lower bound $\Omega(\sqrt{n})$ on the quantum query complexity for $f\colon[n]\to Z$ for a finite set $Z$, where $[n]$ denotes the set $\{1,\dots,n\}$. Aaronson and Shi [2] improved the lower bound to $\Omega(n^{2/3})$ by introducing a novel technique to the polynomial method to bound the degrees of multi-variate polynomials that approximate symmetric Boolean functions. Ambainis [6] then provided the breakthrough upper bound $O(n^{2/3})$ by developing a general framework of quantum search based on quantum walk. His work was followed by many generalizations and variants that form one of the mainstreams of quantum algorithms (e.g.,[30, 27, 8, 23]).
@@
As for claw detection for $f\colon[F]\to Z$ and $g\colon[G]\to Z$, where $F$ and $G$ are positive integers such that $F\leq G$ and $Z$ is a finite set, Buhrman et al. [14] proved that the quantum query complexity is $\Omega(\sqrt{G})$ and $O(\sqrt{G}+F^{1/2}G^{1/4})$. Zhang [33] then proved the lower bound $\Omega(\sqrt{G}+(FG)^{1/3})$ on the quantum query complexity by providing the reduction from collision detection with the promise that the input function with domain $[FG]$ is either one-to-one or two-to-one and using the lower bound $\Omega((FG)^{1/3})$ [2] for that promised case. Tani [31] then proved the matching upper bound $O(\sqrt{G}+(FG)^{1/3})$ by running a quantum-walk-based search on a particular graph product of two Johnson graphs corresponding to the two input functions.
@@
These tight bounds depend only on the domain size(s) of the input function(s). This means that these bounds are tight if the function range is sufficiently large. Indeed, the lower bound $\Omega(n^{2/3})$ of element distinctness was initially proved when the range size is $\Omega(n^{2})$ [2]. In some cases of smaller range size, however, it is still non-trivial to detect a collision (e.g., a collision for a given function $f\colon[n]\to[n]$ since $f$ can be one-to-one). Moreover, in some applications, such as cryptographic ones, the range size is often known and can be a relatively small value. Ambainis [4] and Kutin [25] independently proved the same lower bound $\Omega(n^{2/3})$ holds even for the range of size $n$. In particular, Ambainis’s argument is so general that it can be applied to any function $\Phi$ from a set of functions $f$ to $\{0,1\}$ if $\Phi$ is invariant under the permutations $\pi$ and $\sigma$ over the domain and range, respectively, of $f$, i.e., $\Phi(f)=\Phi(\sigma f\pi)$ (e.g., if $\Phi$ tells the existence of a collision for $f$, then it is easy to see that $\Phi(f)=\Phi(\sigma f\pi)$).
@@
As for claw detection, Zhang’s lower bound holds if the range size of input functions is $\Omega(FG)$. This condition seems very strong, considering that it is still non-trivial  to detect a claw for a given pair of functions $f\colon[F]\to[F+G]$ and $g\colon[G]\to[F+G]$ with $F\leq G$, since the images of $f$ and $g$ can be disjoint. Therefore, it is natural to ask whether the lower bound $\Omega(\sqrt{G}+(FG)^{1/3})$ still holds for a smaller range, such as $[F+G]$, and, more ambitiously, whether a multi-function version of Ambainis’s argument exists. From a practical point of view, relaxing the condition on the function range for tight bounds could contribute to more accurate security analysis of post-quantum ciphers, such as those based on supersingular elliptic curve isogenies [22], one form (SIKE) of which had been considered until the final round of the post-quantum standardization at NIST.
@@ 1.2 Polynomial Method
There are two primary lower-bound techniques for quantum query complexity: the polynomial method [9] and the adversary method [3, 19]. Although the polynomial method does not necessarily give optimal lower bounds [5], it often gives non-trivial lower bounds that were (or are) not previously known to be obtainable using other lower-bound techniques (e.g., Refs. [2, 15]). The highlight of the method is the claim that the minimum degree of a polynomial that approximates a Boolean function $\Phi$ is at most twice the number of quantum queries required to compute $\Phi$. This allows us to use the results of polynomial approximation theory to obtain lower bounds on the quantum query complexity of $\Phi$. More concretely, let $f$ be any function from $[n]$ to $[m]$. Then, $f$ can be identified with $nm$ Boolean variables $\{x_{ij}:i\in\left[n\right],j\in\left[m\right]\}$ such that $x_{ij}=1$ if and only if $f\left(i\right)=j$. Let $\Phi$ be a Boolean function that tells whether input $f$ has a certain property (e.g., the property that $f$ has at least one collision). Then, $\Phi$ is a function in the variables $\{x_{ij}\}$. The polynomial method claims that the minimum degree of a polynomial that approximates $\Phi$ within error $\epsilon$ is at most twice the number of quantum queries required to compute $\Phi$ with error probability at most $\epsilon$. For $m\geq n$, suppose that $\Phi^{\prime}$ is the restriction of $\Phi$ to the case where the range of input function is $\left[n\right]$. Then, it is easy to see that if $d$ is the minimum degree of a polynomial that approximates $\Phi$, then the minimum degree of a polynomial that approximates $\Phi^{\prime}$ is at most $d$ (since $\Phi^{\prime}$ is obtained by setting $x_{ij}=0$ for all $i\in[n]$ and $j\in\left[m\right]\setminus\left[n\right]$ in $\Phi$). Ambainis [4] proved that the converse is also true if $\Phi$ is a symmetric property, i.e., a Boolean function invariant under the permutations over domain $[n]$ and range $\left[m\right]$, respectively, of $f$. For instance, following the lower-bound proof of element distinctness by Aaronson and Shi [2], its minimum approximate degree is $\Omega\left(n^{2/3}\right)$ if the range size of input functions is $m\in\Omega(n^{2})$. Thus, the polynomial method together with Ambainis’s argument implies that the quantum query complexity of the element distinctness problem is $\Omega\left(n^{2/3}\right)$ even if the function range is $\left[n\right]$. This relaxes the condition (i.e., $m=\Omega(n^{2})$) on the range size of input functions required in Ref. [2] required to obtain the optimal lower bound.
@@ 1.3 Our Contribution
Our main contribution is to prove a multi-function version of Ambainis’s argument. As an application of our argument to two functions $f\colon[F]\to[M]$ and $g\colon[G]\to[M]$, we show that the optimal quantum lower bound for claw detection proved for range size $FG$ still holds for smaller range size $F+G$. Moreover, using this new lower bound, we obtain a lower bound for an even smaller range $[M]$ with $2\leq\left|M\right|<F+G$. In the following, we summarize our results more concretely. To simplify the notation, we represent by $\mathcal{R}^{\mathcal{D}}$ the set of functions from $\mathcal{D}$ to $\mathcal{R}$ for any ordered sets $\mathcal{D}$ and $\mathcal{R}$. For instance, if $f$ is a function from $[F]$ to $[M]$, then $f\in[M]^{[F]}$.
@@ 1.3.1 Approximate Degree of Multisymmetric Properties
For $F,G,M\in\mathbb{N}$ such that $M\geq F+G$, let $\Phi\colon[M]^{[F]}\times[M]^{[G]}\to\{0,1\}$ be a bisymmetric Boolean function of a pair of functions $f\colon[F]\to[M]$ and $g\colon[G]\to[M]$. In addition, for every $M^{\prime}$ with $F+G\leq M^{\prime}<M$, $\Phi^{\prime}\colon[M^{\prime}]^{[F]}\times[M^{\prime}]^{[G]}\to\{0,1\}$ be the restriction of $\Phi$ to the case where the input functions are $f\colon[F]\to[M^{\prime}]$ and $g\colon[G]\to[M^{\prime}]$. Then, for $0<\epsilon<1/2$, the minimum degree of a polynomial that $\epsilon$-approximates $\Phi$ is equal to the minimum degree of a polynomial that $\epsilon$-approximates $\Phi^{\prime}$.
@@
The definition of bisymmetric properties can be generalized in a straightforward way: Let $\Phi\colon\mathcal{F}_{1}\times\cdots\times\mathcal{F}_{k}\to\{0,1\}$ be a Boolean function over the set of $k$-tuples of functions in $\mathcal{F}_{1}\times\cdots\times\mathcal{F}_{k}$, where $\mathcal{F}_{\ell}\colon[F_{\ell}]\to[M]$ for $F_{\ell},M\in\mathbb{Z}_{+}$. As in the case of two functions, we can see that $\Phi$ depends on Boolean variables $x^{(\ell)}_{ij}\in\{0,1\}$ for $(\ell,i,j)\in[k]\times[F_{\ell}]\times[M]$ such that $x^{(\ell)}_{ij}=1$ if and only if $f_{\ell}\left(i\right)=j$. We say that $\Phi$ is $k$-symmetric, if $\Phi(f_{1},\dots,f_{k})$ is invariant under the permutations over $F_{\ell}]$ for each $\ell\in[k]$, respectively, and the common range $[M]$ of $f_{\ell}$’s, i.e.,
@@
for every permutation $\pi_{F_{i}}$ over $F_{i}]$ for $i\in[k]$ and every permutation $\sigma$ over $[M]$. Then, we have a generalization of Theorem 1.
@@
Let $k\geq 2$. For $F_{\ell}\ (\ell\in[k])$ and $M\in\mathbb{N}$ such that $M\geq F_{1}+\dots+F_{k}$, $\Phi:\mathcal{F}_{1}\times\dots\times\mathcal{F}_{k}\to\{0,1\}$ be a $k$-symmetric Boolean function over the set of $k$-tuples of functions $(f_{1},\dots,f_{k})\in\mathcal{F}_{1}\times\dots\times\mathcal{F}_{k}$, where $\mathcal{F}_{\ell}$ denotes $M]^{[F_{\ell}]}$ for each $\ell$. In addition, for every $M^{\prime}$} such that $F_{1}+\dots+F_{k}\leq{M^{\prime}}<M$, let $\Phi^{\prime}:\mathcal{F}^{\prime}_{1}\times\dots\times\mathcal{F}^{\prime}_{k}\to\{0,1\}$ be the restriction of $\Phi$ to the domain ${M^{\prime}}]^{[F_{1}]}\times\dots\times[{M^{\prime}}]^{[F_{k}]}$, i.e., the $k$-symmetric Boolean function over the set of $k$-tuples of functions $(f^{\prime}_{1},\dots,f^{\prime}_{k})\in\mathcal{F}^{\prime}_{1}\times\dots\times\mathcal{F}^{\prime}_{k}$, where $\mathcal{F}^{\prime}_{\ell}$ denotes ${M^{\prime}}]^{[F_{\ell}]}$ for each $\ell\in[k]$, such that $\Phi^{\prime}(f^{\prime}_{1},\dots,f^{\prime}_{k})=\Phi(f_{1},\dots,f_{k})$ for all functions $f^{\prime}_{1},\dots,f^{\prime}_{k},f_{1},\dots,f_{k}$ satisfying $f^{\prime}_{\ell}(x)=f_{\ell}(x)$ for all $x\in[F_{\ell}]$ and all $\ell\in[k]$.
@@
With the promise that the union of the images of input functions, i.e., $\bigcup_{\ell\in[k]}{f_{\ell}([F_{\ell}])}$, is upper-bounded in size by $\mathfrak{U}$, Theorem 2 also holds for every $M^{\prime}$ with $\mathfrak{U}\leq M^{\prime}<M$ with a slight modification to the proof. Slightly general versions of the above theorems are Theorems 12 and 15 proved in Sec.   3.